a drink mix label sates the 2 tablespoons of mix should be added to every cps of water. if a container has 40 cups of wateer in it, how much mix should be added to the water? you are supposed to use a porportain, show workthanks so much!.
Answer by josgarithmetic(39613)

Keep the units with their corresponding number in the parts of the ratio. You can use (2 tablspoonsOfMix)/(1 CupsOfWater), or you can use (1 CupsOfWater)/(2 tablespoonsOfMix).

The ratio you use which contains the unknown part must match the form of the known ratio that you choose. You have as given, 40 cups of water, and you need to find how many tablespoons of drink mix.

You can use EITHER of these:
2/1=q/40
OR
1/2=40/q

Find q.